f a chord of the parabola y^2 = 4ax is a normal at one of its ends, show that its mid-point lies on the curve 2(x − 2a) = y^2/a + 8a^3/y^2
1
Expert's answer
2020-07-01T18:31:58-0400

Equation of the normal chord at any point (at2,2at) of the parabola is

y+tx=2at+at3 .....(1)

Equation of the chord with midponit (x1,x22) is T=S1

or yy1-2a(x+x1)=y12-4ax1 or

yy1-2ax=y12-2ax1 ....(2)

since equation (1)and(2) are identical

1/y1 = t/(-2a) = (2at+at3)/t = 2a + {(-2a)/y1}2

or -(y12)/2a+x1 = 2a+ 4a3/(y12)

or x1-2a = (y12)/2a +4a3/(y12)

hence the locus of the middle point (x1,y1) is

x-2a = y2/2a + 4a3/y2

multiply 2 on both sides

2(x-2a) = 2y2/2a + 2(4a3/y2)

2(x-2a) = y2/a + 8a3/y2

hence midpoint lies on this curve